# Confounder-Free Continual Learning via Recursive Feature Normalization

### v1

@@ 3 Methodology
where ${X}_{i,:}$ is the $i^{\text{th}}$ row of ${X}$. If we represent $R(N)=\sum_{i=1}^{N}{X}_{i,:}{X}_{i,:}^{\top}$ and $Q(N)=\sum_{i=1}^{N}{z}_{i}{X}_{i,:}$, this is equivalent to writing $\beta=R(N)^{-1}Q(N)$.
@@ 3.4 Regularization
Table 1: Quantifying metrics for the synthetic dataset in continual learning. ACCd, BWTd (×$\times 10^{-2}$), and FWTd mean and standard deviation for different methods and datasets. A total of 5 runs were performed with different model initialization seeds. We compare against prior works—BR-Net (Adeli et al., 2020a), MDN (Lu et al., 2021), and P-MDN (Vento et al., 2022). MDN, a CNN was trained separately on each training stage and then evaluated against all stages. Best results in bold, second-to-best, underlined.
@@ 4 Experimental Results
[…]
@@ 4.4 Static Learning
ABCD Sex Classification: We used T1w MRIs from the ABCD (Adolescent Brain Cognitive Development) study (Casey et al., 2018) for the task of binary sex classification. Within a cross-sectional study setting, we take 10686 baseline (i.e., first visit) MRIs, confounded by scores from the Pubertal Development Scale (PDS)—a validated measure of pubertal stage identified through self-assessment. PDS is a confounder because it is significantly (Suppl. E) larger in girls (2.175 $\pm$ 0.9) than in boys (1.367 $\pm$ 0.6) (Adeli et al., 2020b). PDS categorizes participants as either (1) pre-pubertal, (2) early-pubertal, (3) mid-pubertal, (4) late-pubertal, or (5) post-pubertal (Carskadon & Acebo, 1993).
@@ Appendix C Synthetic Dataset for Static Learning
dcor2 (group 2) 0.005 $\pm$ 0.001 0.001 $\pm$ 0.000 $0$$\pm$ 0.000 0.001 $\pm$ 0.000 0.008 $\pm$ 0.017
We construct this dataset, following Adeli et al. (2020a) and Lu et al. (2021), by generating 2048 images of size 32$\times$32, equally divided between two groups (categories). Each image consists of 3 Gaussian kernels: two on the main diagonal, i.e., quadrants 2 and 4, whose magnitudes are controlled by parameter $\sigma_{A}$, and one on the off-diagonal, i.e., quadrant 3, whose magnitude is controlled by $\sigma_{B}$ (see Figure 7). Differences in the distributions of $\sigma_{A}$ between the two groups are associated with the main effects (true discrimination cues) that should be learned by the system, whereas $\sigma_{B}$ is a confounding variable. An unbiased system will only use information from the main effects for categorization. Both $\sigma_{A}$ and $\sigma_{B}$ are sampled from the distribution $\mathcal{U}(1,4)$ for group 1, and $\mathcal{U}(3,6)$ for group 2. Since there is an overlap in the sampling range of the main effects between the two groups, the theoretical maximum accuracy that the system can achieve, were it to not depend on the confounding variable to make discrimination choices, would be $0.8331-\left(\frac{1}{2}\right)\mathcal{P}[\sigma_{A}\in\mathcal{U}(3,4)]=1-\left(\frac{1}{2}\right)\left(\frac{1}{3}\right)=$ 0.833.
@@ Appendix D A Second Synthetic Continual Learning Dataset
While we explore a synthetically generated environment in Section 4.1 to theoretically quantify the influence of confounding variables and the deviation of the model from achieving an un-confounded performance, here we construct a slightly more complex setup wherein we vary two different axes of possibly confounding information: position and intensity. More specifically, we generate 1024 32$\times$32 images that are implicitly broken down into 16 8$\times$8 grids. The top left and bottom right grids contain Gaussian kernels of intensity $\sigma_{A}$, denoting the main effects. The confounder is represented by a Gaussian kernel of intensity $\sigma_{B}$, whose position varies from the bottom left to top right of the image over 4 different training stages. Both $\sigma_{A}$ and $\sigma_{B}$ are sampled from the distribution $\mathcal{U}(3,5)$ for group 1, and $\mathcal{U}(4,6)$ for group 2. A fair classifier should remain unaffected by confounder information, irrespective of its position within the image. Such a setup also allows us to compute theoretical maximum accuracy achievable for each training stage that we can validate methods against. Table 8 presents the results of the experiment, where we can see that R-MDN again allows the model to learn confounder-free features.
@@ E.4 ABCD Study
Table 10: Variable distributions across boys and girls in the ABCD study. Mean and standard deviation for age and pubertal development scale (PDS), and the number of subjects of each race in the study across boys and girls. PDS is an integer between 1-5. Differences are significant across boys and girls for age and PDS (measured using a two-sample t-test) but not race. Girls have a higher PDS than boys in the study. All values are for the first visit of each subject.
@@
Age (in months) 119.17 $\pm$ 7.563 118.81 $\pm$ 7.520 $<$0.001
PDS 1.367 $\pm$ 0.615 2.175 $\pm$ 0.904 $<$0.001
White 5954 5370
Black 1510 1558
Race/ethnicity Hispanic 2186 2084 $>$0.001
Asian 214 232
Other 1162 1090
@@ F.2 ABCD Sex Classification
Raw MRI images were downloaded, skull-stripped, and affinely registered to the MNI 152 template (Mazziotta et al., 1995, 2001a, 2001b). Data augmentation involved removing MRIs for all subjects that did not have an associated PDS score recorded. We downscaled all MRIs to 64$\times$64$\times$64 volumes, performed random one voxel shift and one degree rotation in all three Cartesian directions, and random left-right flip (since sex affects the brain bilaterally (Hill et al., 2014; Hirnstein et al., 2019)) for training images. To evaluate the models, we perform 5 runs of 5-fold cross validation across different model initialization seeds, with images split by subject and site ID, and having approximately an equal number of boys and girls in each fold.
@@ F.4 HAM10000 Skin Lesion Classification
The dataset was first downsampled to 64$\times$64$\times$64 and then divided into five training stages based on age groups: $<20$, $[20,40)$, $[40,50)$, $[50,60)$, and $\geq 60$.
  • Stage 1: 50% of the images came from $<20$, 30% from $[20,40)$, 10% from $[40,50)$, 5% from $[50,60)$, and 5% from $\geq 60$.
  • Stage 2: 5% from $<20$, 50% from $[20,40)$, 30% from $[40,50)$, 10% from $[50,60)$, and 5% from $\geq 60$.
  • Stage 3: 5% from $<20$, 5% from $[20,40)$, 50% from $[40,50)$, 30% from $[50,60)$, and 10% from $\geq 60$.
  • Stage 4: 10% from $<20$, 5% from $[20,40)$, 5% from $[40,50)$, 50% from $[50,60)$, and 30% from $\geq 60$.
  • Stage 5: 30% from $<20$, 10% from $[20,40)$, 5% from $[40,50)$, 5% from $[50,60)$, and 50% from $\geq 60$.
@@ F.5 ADNI Diagnostic Classification
The dataset was first downsampled to 64$\times$64$\times$64 and then divided into five training stages based on acquisition year bins: $\leq$2010, 2011, 2012, 2013, and $\geq$2014.
@@ Appendix G Visualizing Results for Continual and Static Learning
Here we visualize the results we present through tables in Sections 4.1, 4.2, and 4.4.

### v2

@@ 3 Methodology
where ${X}_{i,:}$ is the $i^{\text{th}}$ row of X{X}. If we represent $R(N)=\sum_{i=1}^{N}{X}_{i,:}{X}_{i,:}^{\top}$ and $Q(N)=\sum_{i=1}^{N}{z}_{i}{X}_{i,:}$, this is equivalent to writing $\beta=R(N)^{-1}Q(N)$.
@@ 3.4 Regularization
Table 1: Quantifying metrics for the synthetic dataset in continual learning. ACCd, BWTd ($\times 10^{-2}$), and FWTd mean and standard deviation for different methods and datasets. A total of 5 runs were performed with different model initialization seeds. We compare against prior works—BR-Net (Adeli et al., 2020a), MDN (Lu et al., 2021), and P-MDN (Vento et al., 2022). MDN, a CNN was trained separately on each training stage and then evaluated against all stages. Best results in bold, second-to-best, underlined.
@@ 4 Experimental Results
We train on a continuum of data by slightly modifying the setting described by Lopez-Paz & Ranzato (2017): A 4-tuple $(a_{i},x_{i},y_{i},s_{i})$ for $i\in[N]$, where $a_{i}\in{\bm{A}}$ is the input, $x_{i}\in{\bm{X}}$ is the confounder, $y_{i}\in{\bm{Y}}$ is the label, and $s_{i}\in{\bm{S}}$ is the training stage descriptor, satisfies local iid, i.e., $(a_{i},x_{i},y_{i})\overset{\text{iid}}{\sim}\mathcal{P}_{s_{i}}({\bm{A}},{\bm{X}},{\bm{Y}})$. Our goal is to learn a classifier $g:{\bm{A}}\times{\bm{S}}\rightarrow{\bm{Y}}$, that is able to predict the label $y$ associated with an unseen 2-tuple $(a,s)$, where $(a,y)\sim\mathcal{P}_{s}$ at any point during or after training on the $\bm{S}$} stages, in a way such that it does not use confounder information in $x$.
@@ 4.4 Static Learning
ABCD Sex Classification: We used T1w MRIs from the ABCD (Adolescent Brain Cognitive Development) study (Casey et al., 2018) for the task of binary sex classification. Within a cross-sectional study setting, we take 10686 baseline (i.e., first visit) MRIs, confounded by scores from the Pubertal Development Scale (PDS)—a validated measure of pubertal stage identified through self-assessment. PDS is a confounder because it is significantly (Suppl. E) larger in girls (2.182 $\pm$ 0.9) than in boys (1.372 $\pm$ 0.6) (Adeli et al., 2020b). PDS categorizes participants as either (1) pre-pubertal, (2) early-pubertal, (3) mid-pubertal, (4) late-pubertal, or (5) post-pubertal (Carskadon & Acebo, 1993).
@@ Appendix C Synthetic Dataset for Static Learning
dcor2 (group 2) 0.005 $\pm$ 0.001 0.001 $\pm$ 0.000 0.000 $\pm$ 0.000 0.001 $\pm$ 0.000 0.008 $\pm$ 0.017
We construct this dataset, following Adeli et al. (2020a) and Lu et al. (2021), by generating 2048 images of size $\times32$, equally divided between two groups (categories). Each image consists of 3 Gaussian kernels: two on the main diagonal, i.e., quadrants 2 and 4, whose magnitudes are controlled by parameter $\sigma_{A}$, and one on the off-diagonal, i.e., quadrant 3, whose magnitude is controlled by $\sigma_{B}$ (see Figure 7). Differences in the distributions of $\sigma_{A}$ between the two groups are associated with the main effects (true discrimination cues) that should be learned by the system, whereas $\sigma_{B}$ is a confounding variable. An unbiased system will only use information from the main effects for categorization. Both $\sigma_{A}$ and $\sigma_{B}$ are sampled from the distribution $\mathcal{U}(1,4)$ for group 1, and $\mathcal{U}(3,6)$ for group 2. Since there is an overlap in the sampling range of the main effects between the two groups, the theoretical maximum accuracy that the system can achieve, were it to not depend on the confounding variable to make discrimination choices, would be $1-\left(\frac{1}{2}\right)\mathcal{P}[\sigma_{A}\in\mathcal{U}(3,4)]=1-\left(\frac{1}{2}\right)\left(\frac{1}{3}\right)=0.833$.
@@ Appendix D A Second Synthetic Continual Learning Dataset
While we explore a synthetically generated environment in Section 4.1 to theoretically quantify the influence of confounding variables and the deviation of the model from achieving an un-confounded performance, here we construct a slightly more complex setup wherein we vary two different axes of possibly confounding information: position and intensity. More specifically, we generate 1024 $\times32$ images that are implicitly broken down into 16 $\times8$ grids. The top left and bottom right grids contain Gaussian kernels of intensity $\sigma_{A}$, denoting the main effects. The confounder is represented by a Gaussian kernel of intensity $\sigma_{B}$, whose position varies from the bottom left to top right of the image over 4 different training stages. Both $\sigma_{A}$ and $\sigma_{B}$ are sampled from the distribution $\mathcal{U}(3,5)$ for group 1, and $\mathcal{U}(4,6)$ for group 2. A fair classifier should remain unaffected by confounder information, irrespective of its position within the image. Such a setup also allows us to compute theoretical maximum accuracy achievable for each training stage that we can validate methods against. Table 8 presents the results of the experiment, where we can see that R-MDN again allows the model to learn confounder-free features.
@@ E.4 ABCD Study
Table 10: Variable distributions across boys and girls in the ABCD study. Mean and standard deviation for age and pubertal development scale (PDS), and the number of subjects of each race in the study across boys and girls. PDS is an integer between 1-5. Differences are significant across boys and girls for PDS (measured using a two-sample t-test) but not age or race. Girls have a higher PDS than boys in the study. All values are for the first visit of each subject.
@@
Age (in months) 119.18 $\pm$ 7.562 118.80 $\pm$ 7.525 >>0.001
PDS 1.372 $\pm$ 0.622 2.182 $\pm$ 0.907 <<0.001
White 2953 2659
Black 702 735
Race/ethnicity Hispanic 1060 999 >>0.001
Asian 103 114
Other 567 529
@@ F.2 ABCD Sex Classification
Raw MRI images were downloaded, skull-stripped, and affinely registered to the MNI 152 template (Mazziotta et al., 1995, 2001a, 2001b). Data augmentation involved removing MRIs for all subjects that did not have an associated PDS score recorded. We downscaled all MRIs to $\times64$$\times$64 volumes, performed random one voxel shift and one degree rotation in all three Cartesian directions, and random left-right flip (since sex affects the brain bilaterally (Hill et al., 2014; Hirnstein et al., 2019)) for training images. To evaluate the models, we perform 5 runs of 5-fold cross validation across different model initialization seeds, with images split by subject and site ID, and having approximately an equal number of boys and girls in each fold.
@@ F.4 HAM10000 Skin Lesion Classification
The dataset was first downsampled to $\times64$$\times$64 and then divided into five training stages based on age groups: <20<20, $[20,40)$, $[40,50)$, $[50,60)$, and $\geq 60$.
  • Stage 1: 50% of the images came from <20<20, 30% from $[20,40)$, 10% from $[40,50)$, 5% from $[50,60)$, and 5% from $\geq 60$.
  • Stage 2: 5% from <20<20, 50% from $[20,40)$, 30% from $[40,50)$, 10% from $[50,60)$, and 5% from $\geq 60$.
  • Stage 3: 5% from <20<20, 5% from $[20,40)$, 50% from $[40,50)$, 30% from $[50,60)$, and 10% from $\geq 60$.
  • Stage 4: 10% from <20<20, 5% from $[20,40)$, 5% from $[40,50)$, 50% from $[50,60)$, and 30% from $\geq 60$.
  • Stage 5: 30% from <20<20, 10% from $[20,40)$, 5% from $[40,50)$, 5% from $[50,60)$, and 50% from $\geq 60$.
@@ F.5 ADNI Diagnostic Classification
The dataset was first downsampled to $\times64$$\times$64 and then divided into five training stages based on acquisition year bins: $\leq$2010, 2011, 2012, 2013, and $\geq$2014.
@@ Appendix G Visualizing Results for Continual and Static Learning
Here we visualize the results we present through tables in Sections 4.1, 4.2, and $4.$